# Human Attention During Localization of Memory Bugs in C Programs

### v1

@@ 4.2 RQ2: Visual Attention Patterns
A novel observation of this paper is that people read several functions outside of the one where the bug is located (a mean of 29), though visual attention is concentrated among relatively few. Table VIII shows that programmers spend 75% of their visual effort on 22% of the functions that they view. In rough practical terms, for a typical bug a typical person read 29 functions but spent 75% of their visual effort on 6 of them. Figure 3 shows a representative example of this pattern with the bug firefly: The x-axis shows one bar for each function viewed and the height of the bar on the y-axis is the number of fixations. Six study participants did this task. And while the number of functions viewed varied substantially (min 11, max 91), the pattern is consistent of concentrated visual attention. Put succinctly, around 75% of visual attention is paid to 25% of functions that are read.
@@
We observe that the distribution of visual effort over lines is concentrated to a lesser degree than over functions. Consider Table IX. Around 32% of the lines read occupied 75% of the total number of fixations on code lines. The example in Figure 4 for bug firefly shows this pattern. While some participants focused on a small number of lines (e.g., p10), it was common to spread out one’s visual attention more, as indicated with the longer sections marked in Figures 3 and  4 in blue.
Around 60% of the participants’ fixations were spent on code on average, though this number was highly variable. The remaining fixations mostly spent reading the bug report details (a near-trivial amount of time was spent navigating the IDE). Recall from Figure 1 that participants had code, report, navigation, and AI areas visible and that they were not permitted to leave the IDE. Figure 5 shows the percent of visual attention (measured by fixations) each participant spent on code during each bug. We caution against drawing strong conclusions generalizing participant behavior in terms of region fixated on because we observe a range from 10% to 88% for some bugs. Some participants spend the vast majority of their effort reading code, while others spend that effort reading the bug report.
A final observation is that relatively little outside help was requested, as measured by the number of queries people asked in the AI help window and visual effort spent on this area. Table X shows only a handful of queries were made, with 5 of 11 participants not requesting AI help.
@@
Figure 5: Percent fixations on code for each bug. Each point represents one participant’s percent of time looking at code versus other regions while working on a particular bug. Code corresponds to area B in Figure 1. Note not all participants worked on all bugs.
TABLE X: Participant AI Query Counts and Percentages
Participant AI Queries Percent Gazes at AI Window
p1 0 0.00%
p2 5 7.89%
p3 0 0.00%
p4 6 4.09%
p5 8 7.46%
p6 0 0.18%
p7 0 0.05%
p8 2 2.03%
p9 4 2.09%
p10 2 1.44%
p11 0 0.00%
4.3 RQ3: Attention and Task Outcome
@@
Figure 6: Percent fixations on a “correct bug line” from lowest to highest “where accuracy” score for each participant-task. More fixations on correct locations are observed with higher accuracy, but correct lines were still read even at low accuracy.
@@
Overall we observed differences in terms of regression rate and mean euclidean distance between fixations for higher versus lower performance on tasks. We did not observe a difference for fixation count or duration, or unique lines or methods read. Our findings are conceptually similar to reports from Software Engineering and other research areas that higher or lower performance on tasks can be associated with greater focus on certain thematically-relevant components [71, 72, 73, 74, 75]. This section discusses the nature of these differences in our study and key caveats.
@@
First consider Figure 6 which shows the percent of fixations on a “correct bug line.” A “correct bug line” is a line in the code that would have been marked as a correct answer to the “where accuracy” question (see Section 3.6 for definitions). Each bar in Figure 6 is a task by a participant. The bar height is the percent of fixations during that task that the participant looked at a correct bug line. While the percentages themselves should not be over-interpreted, what we observe is that even during tasks with “where accuracy” of 2 or 3, participants often fixated on correct bug lines at least briefly. This means that participants failed to mark these lines as bug locations even though we have evidence that they read those locations.
Figure 7: As ‘Where Accuracy” increases, the mean distance between fixations tends to decrease, and the regression rate tends to increase.
As an initial point for understanding differences in attention based on task performance, we calculated six metrics of visual attention and split them based on high and low performance. We define “high” as “where accuracy” scores 4-5 and “low” as “where accuracy” scores 1-2.
@@
TABLE XI: Statistical summary of eye metrics and accompanying boxplots. Low group means accuracy scores $\leq$2. High group means scores $\geq$4.
@@
fixation count Low 830.7 45.0 2101.0 516.7 0.284
High 982.3 281.0 2411.0 624.8
fixation duration Low 416.0 302.4 502.4 52.5 0.161
High 442.9 333.2 596.4 73.7
regression rate Low 47.5 15.6 67.1 14.6 0.006
High 57.8 21.0 70.5 12.2
mean euclidean dist. Low 141.5 115.5 171.4 14.2 0.003
High 126.9 106.3 148.2 11.5
unique lines read Low 203.7 34.0 392.0 116.4 0.255
High 175.2 36.0 385.0 89.6
unique methods read Low 30.9 5.0 126.0 30.0 0.320
High 22.3 3.0 58.0 15.7
Figure 8: Distribution of bugs per participant grouped by “Where Accuracy.” Accuracy level 2 was most common, but participants were in all groups. Asterisk shows three rounded accuracy scores.
@@
To preserve a conservative analysis, we use five metrics from earlier in this paper: fixation count and duration, regression rate, unique lines and methods read. One additional metric is Mean Euclidean Distance, the mean distance from the x,y coordinates of each fixation to the next fixation. Higher mean distance indicates more eye movement range. We performed a Mann-Whitney U test of low/high scores for each metric. We chose this test to be conservative in not assuming a normal distribution and to be robust to outliers. Table XI shows comparisons for each metric. We observe a statistically-significant difference in regression rate and mean euclidean distance. We observe differences in other metrics but these differences are not statistically-significant. Figure 7 shows regression rate and mean euclidean distance averaged for each “where accuracy” score. The general pattern is that regression rate is higher as “where accuracy” increases, while mean euclidean distance is lower.
@@
Consider the examples in Figures 9-12 to illustrate potential differences in success versus failure cases. In these figures, each bar is a fixation on code (specifically area B indicated in Figure 1). Areas with zero-height bars indicate fixations on other parts of the screen or offscreen. The height of the bar is the duration of the fixation in milliseconds. Colors indicate “correct bug lines” (recall that only one correct bug line is necessary even if multiple are possible). Figure 9 shows a successful case on bug stonefly while Figure 10 shows a failure case. Likewise Figure 11 shows a successful case for bug weevil while Figure 12 shows a failure case.
@@
For bug stonefly, Figure 9 shows the participant reading a correct bug location at line 97 of rdb.c around the seven minute time mark. The participant takes another eleven minutes to read other locations and conclude the session and write a correct answer to the problem. In contrast, Figure 10 shows a different participant reading several areas of code including correct bug lines, but for fewer fixations and less time per fixation – ultimately that participant provided an incorrect answer. A similar pattern is evident in Figure 11 where a participant had several fixations on line 1222 of http_client.c around the 18 minute mark, then continued to read other areas of code, before regressing to that location again and writing it as an answer at the end of the session. And the failure case in Figure 12 again shows fixations on the correct location early in the session but for less time, with the participant ultimately providing an incorrect answer.
@@
A caveat to these findings is summarized in Figure 8: the low and high outcome groups are not balanced by participant because some participants performed less well overall, so some portion of our measurement could be of participant skill versus globally-relevant signs of success or failure. Thus we confine ourselves to conclusions that: 1) we observe higher regression rate and lower mean fixation distance among successful bug location attempts in our study, 2) success or failure is not necessarily explained by having seen a correct bug line but may be better explained by regression rate, and 3) participants did not necessarily report an answer immediately after reading a correct bug line and instead read and reread other areas of code. We view these findings as evidence that visual attention differs during success or failure as a motivation for future work.
Figure 9: Participant p7, success at bug stonefly.
Figure 10: Participant p3, failure at bug stonefly.
5 Discussion / Conclusion
@@
  1. 1.
    We perform a study of human visual attention during bug localization, specifically for memory bugs in C programs. To our knowledge, this study is the first of human visual attention during this particular problem.
  2. 2.
@@
  4. 4.
    We release all data including raw eye tracking observations, post-processed token-level attention, etc., for future researchers.
This work has implications to other areas. Researchers in context-aware neural models and “agentic” AI may be informed by our findings of the extent of code context that humans need to solve localization of memory bugs (RQ2), as a guide to what inputs are necessary for the approaches they design. Researchers in accessibility tools for visual impairments may also find these findings useful to guide what data should be prioritized. Likewise, designers of educational tools may find clues in how to detect when interventions may be needed from our findings and data on attention differences in successful or failure cases. Finally, we present new knowledge of bug localization processes in C, which is of scholarly and public interest.
Figure 11: Participant p11, success at bug weevil.
Figure 12: Participant p5, failure at bug weevil.

### v2

@@ 4.2. RQ2: Visual Attention Patterns
In rough practical terms, for a typical bug a typical person read 23 functions but spent 75% of their visual effort on 5–6 of them. Figure 5(a) shows a representative example of this pattern with the bug firefly: The x-axis shows one bar for each function viewed and the height of the bar on the y-axis is the number of fixations. Eight study participants did this task. And while the number of functions viewed varied substantially (min 9, max 91), the pattern is consistent of concentrated visual attention. Put succinctly, around 75% of visual attention is paid to 25% of functions that are read.
@@
We observe that the distribution of visual effort over lines is concentrated to a lesser degree than over functions. Consider Table 9. Around 31% of the lines read occupied 75% of the total number of fixations on code lines. The example in Figure 5(b) for bug firefly shows this pattern. While some participants focused on a small number of lines (e.g., p10), it was common to spread out one’s visual attention more, as indicated with the longer sections marked in Figures 5(a) and  5(b) in blue.
@@
Figure 6. Percent fixations on code for each bug. Each point represents one participant’s percent of time looking at code versus other regions while working on a particular bug. Code corresponds to area B in Figure 2. Note not all participants worked on all bugs.:
Around 58% of the participants’ fixations were spent on code on average, though this number was highly variable. The remaining fixations mostly spent reading the bug report details (a near-trivial amount of time was spent navigating the IDE). Recall from Figure 2 that participants had code, report, navigation, and AI areas visible and that they were not permitted to leave the IDE. Figure 6 shows the percent of visual attention (measured by fixations) each participant spent on code during each bug. We caution against drawing strong conclusions generalizing participant behavior in terms of region fixated on because we observe a range from 10% to 90% for some bugs. Some participants spend the vast majority of their effort reading code, while others spend that effort reading the bug report.
Table 10. Participant AI Query Counts and Percentage of Gazes on Each Screen Region. Code and Report are combined because participants often opened code and reports in the same region. See Figure 6 for a breakdown between code and report fixations. “Other” includes the iTrace window, out of bounds gazes, and gazes without coordinates. “UNK” indicates where participants changed what was displayed in the region frequently. For example, participant 4 frequently flipped back and forth between the file explorer and code in the same screen region.
Partic.
Queries
to AI
Percent Gazes at
AI Window (D)
Percent Gazes at
Code/Report (A/B)
Percent Gazes at
File Explorer (C)
Percent Gazes at
“Other”
p1 0 0.00% 74.60% 14.40% 11.00%
p2 5 7.89% 78.55% 6.74% 6.83%
p3 0 0.00% 76.43% 9.69% 13.88%
p4 6 4.09% UNK UNK 9.58%
p5 8 7.46% 73.62% 10.94% 7.99%
p6 0 0.18% 78.99% 15.71% 5.12%
p7 0 0.05% 39.89% 2.55% 57.51%
p8 2 1.85% 84.24% 3.45% 10.47%
p9 4 2.09% 76.35% 10.47% 11.09%
p10 2 1.44% 78.82% 6.38% 13.36%
p11 0 0.00% UNK UNK 6.06%
p12 2 2.72% UNK UNK 35.92%
p13 11 12.42% 78.41% 5.08% 4.10%
p14 3 3.97% UNK UNK 18.41%
p15 8 14.51% UNK UNK 8.33%
p16 4 UNK UNK 4.41% 15.08%
p17 6 3.33% 85.94% 6.30% 4.43%
p18 7 1.01% UNK UNK 46.83%
p19 4 2.37% 73.43% 7.46% 16.75%
p20 0 0.39% UNK UNK 6.07%
p21 2 1.64% 80.34% 11.59% 6.43%
A final observation is that relatively little outside help was requested, as measured by the number of queries people asked in the AI help window and visual effort spent on this area. Table 10 shows only a handful of queries were made per participant, with 6 of 21 participants not requesting AI help. Using the AI was not correlated with confidence, accuracy, or difficulty. As instructed, participants did not ask the AI to solve/find the bug or copy and paste significant snippets of code. Instead they asked questions related to how to use calloc (P4, P5, P9, P13, P14, P15, P17, P18, P19) and what is zmalloc (P4, P5, P13, P18, P21), for example.
@@
Figure 7. Percent fixations on a “correct bug line” from lowest to highest “where accuracy” score for each participant-task. More fixations on correct locations are observed with higher accuracy, but correct lines were still read even at low accuracy.:
4.3. RQ3: Attention and Task Outcome
@@
Overall we observed differences in terms of regression rate and mean euclidean distance between fixations for higher versus lower performance on tasks. Although we did not find statistically significant differences in fixation count, fixation duration, or the number of unique lines or methods read, a lack of statistical significance does not imply the absence of an underlying effect; it only means that the current data do not provide sufficient evidence to detect one. Our findings are conceptually similar to reports from Software Engineering and other research areas that higher or lower performance on tasks can be associated with greater focus on certain thematically-relevant components (Carmichael et al., 2010; Eivazi and Bednarik, 2011; Fritz et al., 2014; Jin et al., 2021; Steichen et al., 2014). This section discusses the nature of these differences in our study and key caveats.
@@
First consider Figure 7 which shows the percent of fixations on a “correct bug line.” A “correct bug line” is a line in the code that would have been marked as a correct answer to the “where accuracy” question (see Section 3.6 for definitions). Each bar in Figure 7 is a task by a participant. The bar height is the percent of fixations during that task that the participant looked at a correct bug line.
Key Finding 2: While the percentages themselves should not be over-interpreted, what we observe is that even during tasks with a low “where accuracy”, participants often fixated on correct bug lines at least briefly. This means that participants failed to mark these lines as bug locations even though we have evidence that they read those locations.
@@
Table 11. Statistical summary of eye metrics and accompanying boxplots. Low group means accuracy scores $\leq 3$. High group means accuracy scores $\geq 4$.
@@
fixation count Low 820.4 45.0 2101.0 493.5 0.392
High 888.1 273.0 2411.0 540.2
fixation duration Low 437.1 302.4 792.8 52.5 0.495
High 432.3 316.2 596.4 73.7
regression rate Low 50.7 15.6 77.7 14.6 0.003
High 58.6 21.0 78.9 12.7
mean euclidean dist. Low 136.8 109.9 196.0 15.1 0.0004
High 126.4 106.3 148.2 9.8
unique lines read Low 182.5 31.0 433.0 108.8 0.184
High 162.0 36.0 468.0 99.0
unique methods read Low 26.1 3.0 126.0 22.4 0.068
High 19.2 2.0 71.0 15.8
Figure 8. As ‘Where Accuracy” increases, the mean distance between fixations tends to decrease, and the regression rate tends to increase.:
Figure 9. Boxplots to accompany statistical summary of eye metrics. Low group means accuracy scores $\leq$3. High group means scores $\geq$4.:
As an initial point for understanding differences in attention based on task performance, we calculated six metrics of visual attention and split them based on high and low performance. We define “high” as “where accuracy” scores 4–5 and “low” as “where accuracy” scores 1–3. We chose to group the scores in this way for three reasons. First, for the tasks where participants scored a 3, we looked at the rest of their tasks and determined that most participants who scored a 3 on one task also had more lower scores (1s and 2s) than higher scores (4s and 5s). Second, looking at the regression rate and euclidean distance averages used to make Figure 8, the averages for the “where accuracy” score of 3 was much closer to the averages for the “where accuracy” score of 2 than to the averages for the “where accuracy” score of 4, indicating that the visual attention patterns of 3s are closer to the visual attention patterns of 2s than 4s. Third, in order to receive a score of 3 for “where accuracy”, participants needed to have a “limited idea where to look” according to the rubric, and we do not deem that to be classified as “high performance.” Notably, if we choose to group the scores in different ways, (1s and 2s) versus (3s, 4s, and 5s) or (1s and 2s) versus (4s and 5s), the key findings and statically significant results do not change.
@@
To preserve a conservative analysis, we use five metrics from earlier in this paper: fixation count and duration, regression rate, unique lines and methods read. One additional metric is Mean Euclidean Distance, the mean distance from the x,y coordinates of each fixation to the next fixation. Higher mean distance indicates more eye movement range. We performed a Mann-Whitney U test of low/high scores for each metric. We chose this test to be conservative in not assuming a normal distribution and to be robust to outliers. Table 11 and Figure 9 show comparisons for each metric.
Key Finding 3: We observe a statistically-significant difference in regression rate and mean euclidean distance between bug localizations with low “where accuracy” and high “where accuracy”.
We observe differences in other metrics but these differences are not statistically-significant. Figure 8 shows regression rate and mean euclidean distance averaged for each “where accuracy” score. The general pattern is that regression rate is higher as “where accuracy” increases, while mean euclidean distance is lower.
(a) Participant p7, success at bug stonefly.
(b) Participant p3, failure at bug stonefly.
Figure 10. Comparing successful and failure for cases for bug stonefly
(a) Participant p11, success at bug weevil.
(b) Participant p5, failure at bug weevil.
Figure 11. Comparing successful and failure for cases for bug weevil:
@@
Consider the examples in Figures 10(a)11(b) to illustrate potential differences in success versus failure cases. In these figures, each bar is a fixation on code (specifically area B indicated in Figure 2). Asterisks indicate truncated fixation durations to improve readability. Areas with zero-height bars indicate fixations on other parts of the screen or offscreen. The height of the bar is the duration of the fixation in milliseconds. Colors indicate “correct bug lines” (recall that only one correct bug line is necessary even if multiple are possible). Figure 10(a) shows a successful case on bug stonefly while Figure 10(b) shows a failure case. Likewise Figure 11(a) shows a successful case for bug weevil while Figure 11(b) shows a failure case.
@@
For bug stonefly, Figure 10(a) shows the participant reading a correct bug location at line 97 of rdb.c around the seven minute time mark. The participant takes another eleven minutes to read other locations and conclude the session and write a correct answer to the problem. In contrast, Figure 10(b) shows a different participant reading several areas of code including correct bug lines, but for fewer fixations and less time per fixation – ultimately that participant provided an incorrect answer. A similar pattern is evident in Figure 11(a) where a participant had several fixations on line 1222 of http_client.c around the 18 minute mark, then continued to read other areas of code, before regressing to that location again and writing it as an answer at the end of the session. And the failure case in Figure 11(b) again shows fixations on the correct location early in the session but for less time, with the participant ultimately providing an incorrect answer.
Key Finding 4: In success cases, participants often spend more time on the buggy line or regress back to the buggy line several times while in failure cases, participants often see the buggy line, but do not register it as the answer.
Figure 12. Distribution of bugs per participant grouped by “Where Accuracy.” Accuracy level 2 was most common, but participants were in all groups. Asterisk shows two rounded accuracy scores.:
A caveat to these findings is summarized in Figure 12: the low and high outcome groups are not balanced by participant because some participants performed less well overall, so some portion of our measurement could be of participant skill versus globally-relevant signs of success or failure.
@@
Thus we confine ourselves to conclusions that: 1) we observe higher regression rate and lower mean fixation distance among successful bug location attempts in our study, 2) success or failure is not necessarily explained by having seen a correct bug line but may be better explained by regression rate, and 3) participants did not necessarily report an answer immediately after reading a correct bug line and instead read and reread other areas of code.
We view these findings as evidence that visual attention differs during success or failure as a motivation for future work.
5. Discussion / Conclusion
5.1. Contributions
@@
  1. (1)
    We perform a study of human visual attention during bug localization, specifically for memory bugs in C programs. To our knowledge, this is the first study to examine human visual attention during memory bug localization in C programs.
  2. (2)
@@
  4. (4)
    We release all data including raw eye tracking observations, post-processed token-level attention, etc. for future researchers.
5.2. Intellectual Merit
The intellectual merit/novelty of this paper is that it is a study of C memory bugs. A significant majority of both program comprehension and debugging eye tracking studies utilize Java programs (Obaidellah et al., 2018; Grabinger et al., 2024). The debugging eye tracking studies that use C programs do not focus on memory bugs (Li et al., 2021; Zeng et al., 2022; Sharif et al., 2012; Hou et al., 2013). Conducting a study specifically on C memory bugs is important because memory bugs are one of the most common programming errors in C programs (Van der Veen et al., 2012; The Chromium Project, [n. d.]). Furthermore, C is in the top 4 most popular programming languages (He et al., 2025). To our knowledge, this first study to examine human visual attention during memory bug localization in C programs provides information about a non-studied specific type of problem, allowing for further comparing and contrasting with other C debugging studies and well-studied problems like Java method comprehension.
5.3. Implications
4.2 Programmers spend 75% of their visual effort on 25% of the functions that they view. While the fact that some functions are more important than others when debugging is intuitive, quantifying how many of the functions looked at are the most important provides information that can be used in context-aware neural models and “agentic” AI. For example, an “agentic” AI may determine a number of functions that may be related to a bug, rate them by confidence, and then investigate the top 25% before giving a preliminary answer. Additionally, these numbers provide guidance on the amount of code context that humans require to localize a bug, and therefore may inform how much context should be provided to a neural model to reduce tradeoffs. Researchers in accessibility tools for visual impairments may also find this finding useful to determine how much data should be prioritized.
4.3 Even during tasks with “where accuracy” of 1, 2 or 3, participants often fixated on correct bug lines at least briefly. This finding provides fodder for a future study that investigates why participants dismiss the correct buggy line even after fixating on it.
4.3 As “Where Accuracy” increases, the mean euclidean distance tends to decrease, and the regression rate tends to increase. 4.3 In success cases, participants often spend more time on the buggy line or regress back to the buggy line several times. These differences between success and failure cases may aid designers of educational tools or AI tools. They may use these findings to inform how to detect when interventions may be needed based on eye tracking data. This may help provide targeted instruction to students or trigger an additional training or code review in professional settings.